# Beyond Belief Propagation: Cluster-Corrected Tensor Network Contraction with Exponential Convergence

### v1

@@ I Introduction
We achieve our results through a novel approach that overcomes fundamental limitations of existing methods. In Ref. [50, 51], the authors take inspirations from earlier work in statistical inference [52, 53, 54, 55, 56, 57, 58, 59] to construct a Taylor series known as the loop expansion in order to correct the BP approximation value toward the ground truth. This loop expansion starts from the BP value $Z_{0}$ at zeroth order and sums corrections called loop tensors that grow larger at higher orders. While this seems like a natural tuning knob—–when fully summed, the loop expansion yields ZZ—–the method suffers from a critical flaw: it does not generally converge, even when individual loop tensors decay exponentially. The fundamental problem is that the expansion includes disconnected loops whose number grows combinatorially with size, overwhelming any exponential decay and causing divergence.
@@ II.2 Belief propagation
where rank-one $\triangleright$ denote message tensors and $T_{v}$ is the local tensor at vertex $v$. The messages are normalized to unit norm, $\mu_{v\to w}\star\mu_{v\to w}=1$.
For each $v\in V$, the local contribution to the partition function is then given as,
$Z^{(v)}:=\left[\underset{n\in\mathcal{N}(v)}{\bigotimes}\mu_{n\to v}\right]\star T_{v}$ (4)
The BP vacuum solution to the partition function and the free energy are then given as,
@@
This is graphically shown below. $Z_{0}$ is a approximation to $\mathcal{Z}$.
@@ II.3 Loop series expansion
$\mathbbm{1}=|\mu_{r\to s}\rangle\!\rangle\langle\!\langle\mu_{s\to r}|+\mathcal{P}_{rs}^{\perp}$ (7)
This ensures that $\langle\!\langle\mu_{s\to r}|\mathcal{P}_{rs}^{\perp}|\mu_{r\to s}\rangle\!\rangle=0$, hence $\mathcal{P}_{rs}^{\perp}$ carries contributions orthogonal to the BP vacuum. This is shown in Fig. 1(b), and we ensure normalization of messages, with $\langle\!\langle\mu_{r\to s}|\mu_{r\to s}\rangle\!\rangle=1$ for all $(r,s)\in E$.
@@ III.3 Toy example
III.4 Comparison to previous results
We now compare the cluster expansion to previous approaches. In Ref. [50], the authors introduced the loop series expansion for tensor networks and proposed several strategies to mitigate its combinatorial growth. The cluster expansion can be viewed as a formalization of the intuition underlying the “single-excitation” and “multi-excitation” approximation discussed in Ref. [50]. Our results advance this line of work in three key respects. First, the cluster expansion yields explicit expressions for correction terms at all orders. Second, it applies beyond the thermodynamic limit, whereas the previous methods are restricted to that setting. Lastly, we resolve an open question by showing that sufficiently fast loop decay guarantees exponential convergence of the cluster expansion.
@@ IV.2 Message Passing and Normalization
Message passing is typically efficient, as belief propagation often converges rapidly. However, tensor networks with sign structures may either fail to converge or converge slowly. Additionally, message passing can admit multiple fixed points, some of which may yield poor approximations.
@@ V.2 Cluster Expansion
We now present the central numerical evidence of this work: a systematic comparison between our cluster expansion method and the ‘traditional’ loop series expansion, revealing fundamental differences in how the algorithms scale. Figure 4(d) compares the approximation errors of cluster corrections (dashed, blue) and loop corrections (dotted, red) as functions of correction weight $w\in\{4,6,8,10\}$ for multiple system sizes $L\in\{10,20,40\}$.
@@ VI Discussion
Our work opens several avenues for future research. First, BP is widely used in decoding classical and quantum error-correcting codes [66, 67, 31]. It would be interesting to explore the application of our method to improve the performance of BP-based decoders and provide rigorous error estimates. In particular, it is known that naive BP fails to give a threshold in quantum-LDPC codes because of the degeneracy problem [43]. Since the cluster expansion captures the short-range correlations omitted by the BP approximation, it may help restore a threshold, if not simply improving the decoding performance, which we investigate in a forthcoming work [68].
@@
VII Acknowledgements
We thank Hongkun Chen, Grace Sommers, Rhine Samajdar, David Huse, Sarang Gopalakrishnan, Dima Abanin, Joseph Tindall, Will Staples, and Dries Sels for helpful discussions and collaborations on related projects. Y.F.Z acknowledges support from NSF QuSEC-TAQS OSI 2326767. Parhey’s company is also acknowledged.

### v2

@@ I Introduction
We achieve our results through a novel approach that overcomes fundamental limitations of existing methods. In Ref. [50, 51], the authors take inspirations from earlier work in statistical inference [52, 53, 54, 55, 56, 57, 58, 59] to construct a Taylor series known as the loop expansion in order to correct the BP approximation value toward the ground truth. The naïve loop expansion starts from the BP value $Z_{0}$ at zeroth order and sums corrections called loop tensors that grow larger at higher orders. While this seems like a natural tuning knob—–when fully summed, the loop expansion yields ZZ—–the method suffers from a critical flaw: it does not generally converge, even when individual loop tensors decay exponentially. The fundamental problem is that the expansion includes disconnected loops whose number grows combinatorially with size, overwhelming any exponential decay and causing divergence. The work [50] made the first step in mitigating the combinatorial growth through multiple self-consistent methods.
@@
I.1 Comparison with existing and concurrent work
We compare our results with earlier and concurrent works. Ref. [52, 53] first introduce the loop series expansion in probabilistic graphical models. Ref. [50] generalizes the loop series expansion to tensor networks. Notably, they propose several methods to overcome the divergence of the naïve loop expansion. Their key intuition is the extensivity of free energy which also motivates our paper. In particular, the combinatorial toolbox we will use formalizes the counting argument underlying their “single-excitation” and “multi-excitation” approximation. Our results advance this line of work in three key respects. First, the cluster expansion yields explicit expressions for correction terms at all orders. Second, it applies beyond the thermodynamic limit, whereas the previous methods are restricted to that setting. Lastly, we resolve an open question by showing that sufficiently fast loop decay guarantees exponential convergence of the cluster expansion.
There are also earlier attempts to derive modified series expansion with better convergence than the naïve loop expansion. Notable examples include the linked cluster expansion [60], also known as the cluster-cumulant expansion [61] in the statistical inference literature. Ref. [51] employs the linked cluster expansion as one of the subroutines to compute observables in time-evolved quantum systems. Concurrent to our work, Ref. [62] appeared which employs the linked cluster expansion to perform extensive numerical simulations of quantum ground states. We note that while the linked cluster expansion is conceptually very similar to the cluster expansion derived in our work, they are formally two different series. See Section III.3 for an example. However, we believe that our theory of convergence should apply to linked cluster expansion as well, since both series reflect fundamentally the same physics.
I.2 Organization
@@ II.2 Belief propagation
where rank-one $\triangleright$ denote message tensors and $T_{v}$ is the local tensor at vertex $v$.
To evaluate the BP approximation, we first compute $I_{vw}=\mu_{v\to w}\star\mu_{w\to v}$, that is the inner product between $\mu_{v\to w}$ and $\mu_{w\to v}$ 111note that no complex conjugation is take here, in other words $I_{vw}=I_{wv}$. For each $v\in V$, the local contribution to the partition function is then given as,
$Z^{(v)}:=\left[\underset{n\in\mathcal{N}(v)}{\bigotimes}\frac{\mu_{n\to v}}{\sqrt{I_{vw}}}\right]\star T_{v},$ (4)
where we take the principle branch of the square root when $I_{vw}$ is complex. The BP vacuum solution to the partition function and the free energy are then given as,
@@
This is graphically shown below. $Z_{0}$ is the BP approximation to $\mathcal{Z}$.
@@ II.3 Loop series expansion
$\mathbbm{1}=\frac{|\mu_{r\to s}\rangle\!\rangle\langle\!\langle\mu_{s\to r}|}{I_{rs}}+\mathcal{P}_{rs}^{\perp}$ (7)
This ensures that $\mathcal{P}_{rs}^{\perp}|\mu_{r\to s}\rangle\!\rangle=0$ and $\langle\!\langle\mu_{s\to r}|\mathcal{P}_{rs}^{\perp}=0$, hence $\mathcal{P}_{rs}^{\perp}$ carries contributions orthogonal to the BP vacuum. This is shown in Fig. 1(b),
@@ III.3 Toy example
Finally, this toy example also illustrates the difference between the linked cluster expansion and our cluster expansion. The linked cluster expansion produces $\log(1+Z_{l})$ in one iteration [61], whereas our cluster expansion is an infinite series expansion.
@@ IV.2 Message Passing and Normalization
For numerical stability, we normalize the message vector by its two-norm after each update. This does not affect the result in the end as any constant factor will be cancelled out when dividing by $I_{vw}$. Message passing is typically efficient, as belief propagation often converges rapidly. However, tensor networks with sign structures may either fail to converge or converge slowly. Additionally, message passing can admit multiple fixed points, some of which may yield poor approximations.
@@ V.2 Cluster Expansion
We now present the central numerical evidence of this work: a systematic comparison between our cluster expansion method and the ‘traditional’ loop series expansion, revealing fundamental differences in how the algorithms scale. Figure 4(d) compares the approximation errors of cluster corrections (dashed, blue) and loop corrections (dotted, red) as functions of correction weight $w\in\{4,6,8,10\}$ for multiple system sizes $L\in\{10,20,30\}$.
@@ VI Discussion
While the cluster expansion does not converge in adversarial cases (otherwise there will be severe complexity-theoretic consequences), we believe that it should converge in many physically-relevant setups. In particular, the convergence of the cluster expansion reflects that the system is short-range correlated. Therefore, incorporating small clusters would already account for most of the correlations. This idea is not new in the literature; many earlier works have exploited short-ranged correlations to simply tensor network contractions, for example see [70, 71]. An important theoretical question is to understand what kind of tensor networks admit a good BP approximation with a convergent cluster expansion.
Our work opens several avenues for future applications. First, BP is widely used in decoding classical and quantum error-correcting codes [72, 73, 31]. It would be interesting to explore the application of our method to improve the performance of BP-based decoders and provide rigorous error estimates. In particular, it is known that naive BP fails to give a threshold in quantum-LDPC codes because of the degeneracy problem [43]. Since the cluster expansion captures the short-range correlations omitted by the BP approximation, it may help restore a threshold, if not simply improving the decoding performance, which we investigate in a forthcoming work [74].
@@
Acknowledgements.
We thank Hongkun Chen, Rhine Samajdar, David Huse, Sarang Gopalakrishnan, Dmitry Abanin, Joseph Tindall, Will Staples, and Dries Sels for helpful discussions and collaborations on related projects. We are particularly indebted to Grace Sommers for extensive discussions and cross-checking some of our numerical results. Y.F.Z acknowledges support from NSF QuSEC-TAQS OSI 2326767. The numerical simulations presented in this article were performed on computational resources managed and supported by Princeton University’s Research Computing. Parhey’s company is also acknowledged.
Note added
After the completion of this work, a related work employing cluster expansion on tensor networks was posted [62]. See Section I.1 for detailed discussions.